OpenStudy (dls):

sinx+cosecx=2 sin^8x+cosec^8x=?

OpenStudy (dls):

A better method than binomial expansion? :/

OpenStudy (dls):

i got it :)

OpenStudy (anonymous):

sinx + 1/sinx = 2 sin^2 x +1 = 2sinx sin^2x - 2sinx + 1 = 0 Use Quadratic Formula to find the value of sinx

OpenStudy (dls):

no use of quadratic formula..it is simply (sinx-1)^2=0 sinx=1 ....)

OpenStudy (anonymous):

Any how

OpenStudy (dls):

:)